A spring stores 5 J of energy when stretched by 25 cm. It is kept vertical with the lower end fixed. A block fastened to its other end is made to undergo small oscillations. If the block makes 5 oscillations each second, what is the mass of the block?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. . x = 25cm = 0.25m E = 5J f = 5 So, T = 1/5sec. Now P.E. = (1/2) kx2 ⇒(1/2) kx2 = 5 ⇒ (1/2) k (0.25)2 = 5 ⇒ k = 160 N/m. Again, T = 2π √(m/k) ⇒ 1/5 = 2π √(m/160) ⇒ m = 0.16 kg.